OpenStudy (anonymous):

The following parametric equations trace out a loop. x=3-(2/2)t^2 y=(-2/6)t^3+2t+2 Find the t values at which the curve intersects itself: t = +/- . What is the total area inside the loop? Area = .

OpenStudy (blockcolder):

First, you need to find \(t_1\) and \(t_2\) such that \(x(t_1)=x(t_2)\) and \(t_1\neq t_2\): A little clarification: is \(x= 3-\frac{2}{2}t^2\) or \(x=3-\frac{2}{2t^2}\)?

OpenStudy (blockcolder):

Well, whichever it is, the t-values you're looking for ate \(\large t=\pm\frac{\sqrt{6}}{2}\).

OpenStudy (anonymous):

i think its 2/2, it is very unclear, So I have to equate the x equation?

OpenStudy (blockcolder):

Yeah, so you'll end up with \(3-t_1^2=3-t_2^2 \Rightarrow\) \(t_1^2=t_2^2 \Rightarrow\) \(|t_1|=|t_2| \). WLOG, let \(t_2=-t_1\), since we don't want \(t_1=t_2\). Now, you have to equate \(y(t_1)=y(t_2)\) and use the fact that t2=-t1 to find out the value of t1. Let me know what you got.

OpenStudy (anonymous):

aaahhhhh!!!! I think I got it, after equating both x and finding out that for t_2 to be equal to t_1, one of them has to be the opposite. Then when I equated y(t1)=y(t2) I can just do -(yt2)=(yt2) and then solve for t, getting t=0 and t= sqrt(6). However the one thing I don't understand is how we're allowed to equate both x.